1 ZBol2Harvey Documentation

Predicate Bol++ Z2Bol++ Harvey
¬
∧
∨
⇒
⇔
∀ | •
∃ | •
∃_{1} | •
true
false
Expression

Expression  Bol++ Z2Bol++ Harvey
∈
∉
∩
∪
⊂
⊆
∅
ℙ
\
⊖
↦
×
seq
iseq
=
≠
≤
<
≥
>
−
+
≻
mod
. .
↔
ℕ
ℕ_{1}
ℤ
𝔽
𝔹
↿
↿
⌢
→
\underrightarrow{}
↣
↣
\overrightarrow{→}
\underrightarrow{}
\underrightarrow{↣}
\underrightarrow{}
\begin{matrix} \\ ↣ \end{matrix}
RefName
NumberExpression
SetExpression

2 Z Elements which does not have corresponding elements in Bol

• SymDiff

• Mapsto

• Ffun

• Extract

• Filter

• Definite Description

• Binding Construction

• Unique Existential Quantification

• Tuple Selection

• Schema Projection

• Schema Hiding

• Schema Piping

• Schema Composition

• Schema Piping

3 Changes in Bol grammar

1. add PredicateExpression

4 Obs